DOUBLE-WELL STATES OF UNGERADE SYMMETRY IN H_{2}: FIRST OBSERVATION AND COMPARI-SON WITH AB INITIO CALCULATIONS.

A. DE LANGE, W. HOGERVORST, \underline{W. UBACHS}, Laser Centre, Department of Physics and Astronomy, Vrije Universiteit, Amsterdam, Netherlands; L. WOLNIEWICZ, Nicholas Copernicus University, 87-100 Torun, Poland.

The observation of a new class of long-lived outer well states of ungerade symmetry (B′′B̄^{1}Σ_{u}^{+}) in molecular hydrogen, lying above the ionization threshold, is reported. Rovibrational levels within a potential extended over internuclear separations of R = 7 − 25 a.u. are experimentally investigated in a triple resonance scheme involving a tunable extreme ultra violet (XUV) laser, two tunable infrared lasers, and a fourth laser for ionization. Good agreement (< 0.5 cm^{−1}) with updated ab initio calculations is found for vibrational levels up to v = 26, demonstrating that such calculations can now be extended to this energetic range above ionization, as long as interaction with the Rydberg manifolds is shielded by a barrier. The dynamical behaviour (pre-dissociation and auto-ionization) of this class of ‘u’ symmetry states is remarkably different from similar outer well states of ‘g’ symmetry; this phenomenon can be understood from the structure of doubly-excited electronic states.